A loan of 15,000 is repaid by level repayments at the end of each month for 8 years. The effective annual rate of interest is 5%. Calculate:

a) The monthly repayment.

b) The capital outstanding at the end of the fourth year after the repayment has been made.

c) The interest and capital components of the 49th

instalment.


1
Expert's answer
2022-01-10T17:13:02-0500

effective annual rate of interest:

"r_1=(1+r)^{12}-1=0.05"

 interest rate per period:

"r=1.05^{1\/12}-1=0.0041"


a)

monthly repayment:

"R=\frac{Pr}{1-(1+r)^{-n}}=\frac{15000\cdot0.0041}{1-1.00041^{-96}}=190.31"


b)

capital outstanding at the end of the fourth year:

"BAL=15000(1+0.0041)^{48}-\frac{190.31((1+0.0041)^{48}-1)}{0.0041}=18255.16-10073.00=8182.16"


c)

interest component of the 49th instalment:

"INT=BAL\cdot0.0041=8182.16\cdot0.0041=33.55"


capital component of the 49th instalment:

"PRN=R-INT=190.31-33.55=156.76"